# SeFi-IDE: Semantic-Fidelity Identity Embedding for Personalized Diffusion-Based Generation

### 31 Jan 2024 v1

@@ 3.2 Semantic-Fidelity Token Optimization
Lack of Semantic-Fidelity Control. This problem can be found from two perspectives: (1) Stable Diffusion Model: We observe that even though the face data of celebrities has been included in the training dataset of Stable Diffusion Model, it still fails to achieve perfect semantic-fidelity control for these IDs. For example, “a photo of an old Obama” can not generate corresponding images. (2) Previous Personalized Methods: These works [48, 47] based on Textural Inversion [8] mainly emphasize how to preserve the characteristics of the person and achieve global control over the generated images through text modifications. Although these methods are able to manipulate scenes or styles, they struggle to control fine-grained facial attributes of learned IDs, such as age and expressions.
@@
Disentanglement of $\bm{K}$ and $\bm{V}$ Features. The text prompt will be encoded into text features ($\bm{K}$, $\bm{V}$) by the clip text encoder of Stable Diffusion Model, and then fed into cross-attention layers of U-Nets for conditioning the generated images. Previous methods [19, 35] differentiated the $\bm{K}$ and $\bm{V}$ matrices as position information and object texture features is not appropriate for control facial attributes. As shown in Fig. 5, we found that $\bm{K}$ and $\bm{V}$ features have different semantic meanings. The $\bm{V}$ feature is more ID-related, while the $\bm{K}$ feature is more related to environment factors such as lighting. The disentangled optimization of $\bm{K}$ and $\bm{V}$ features can further improve the ID accuracy and interactive generative ability.
@@
Specifically, the input prompt is fed into the CLIP text encoder, which generates the textual token embedding $\bm{P_{t}}$. The $\bm{P_{t}}$ is fed to a self-attention [38] module, which consists of two self-attention layers with one feed-forward layers. Finally, we evenly divide the 1000 diffusion steps into five stages and optimize one ID embedding as five per-stage tokens $\bm{P}={\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$, while each token consists of a $\bm{V}$ feature and a $\bm{K}$ feature as follows:
$\mathcal{L}_{K,V}=\mathbb{E}_{\bm{z},\bm{t},\bm{P}}[||\epsilon-\epsilon_{\theta}(\bm{z_{t}},\bm{t},(\bm{P_{i}^{K}},\bm{P_{i}^{V}}))||_{2}^{2}],$ (2)
@@
where $\bm{z}$ is the latent code of target image . The $\epsilon$ is the unscaled noise sample, and the $\epsilon_{\theta}$ is the U-Net modules in diffusion Model. The $\bm{t}$ is the optimization step of the diffusion process, and the $\bm{z_{t}}$ is U-Net output of different steps.
@@
Our goal is seamlessly embedding one specific ID into the space of Stable Diffusion Model, which have to achieve accurate ID mapping and fully use the prior from the Stable Diffusion Model to manipulate scenes, actions and facial attributes. Thus, the total optimization objective can be formulated as follows:
@@
Figure 3: Qualitative comparisons with different SOTA methods using different prompts.
4 Experiments
4.1 Experimental Settings
Implementation Details. (1) Pre-Trained Model: In all experiments, we utilize Stable Diffusion 1.4 [29] with default hyper parameters as the pre-trained diffusion-based T2I model. We adopt a frozen CLIP model [31] in the Stable Diffusion Model as the text encoder network. The texts are tokenized into start-token, end-token, and 75 non-text padding tokens. (2) Test Data: The test face images are selected from the CelebA-HQ dataset [15] and then resized to 512. (3) Training Details: For our method, the time for fine-tuning every ID using only one face image is $\sim 5$ minutes ($\sim 1000$ epochs) on one NVIDIA TITAN RTX GPU. We adopt Adam optimizer and set its learning rate as 0.005. The $\lambda$ is set as 0.003.
Baselines. Our task setting is using only a face image to embed ID into the Stable Diffusion Model. Thus, for fair comparisons, we only use a single image for all personalized generation methods but using enough optimization time for different methods. We select five state-of-the-art works as baseline methods for comparisons from two perspectives: (1) Model Fine-Tuning: DreamBooth [30] (learns a unique identifier and fine-tunes the diffusion model to learn from a set of images). Custom Diffusion [19] (retrieves images with similar captions of the target concept and optimizes the cross-attention module with a modifier token), (2) Token Optimization: Textual Inversion [8] (learns a pseudo-word for a concept within a limited number of images for optimization), ProSpect [48] (expands the textual conditioning space with several per-stage textual token embeddings), and Celeb Basis [47] (builds a well-defined face basis module to constrict the face manifold).
Figure 4: The manipulation diversity of our methods, which is shown from stylization, action, and facial attribute.
@@ 4.1 Experimental Settings
Prompt ID Detect ID (Prompt) Prompt ID Quality Train Add (min)
@@
Table 1: Quantitative evaluation between different SOTA methods and ours.
Metrics. Our quantitative evaluation is conducted from objective metrics, user study, parameter amount, and fine-tuning time. (1) Objective Metrics: We select Prompt (CLIP score [31]), ID (ID feature [4] similarity), and Detect (face detection rate [4]). Furthermore, to reflect the essence of T2I generation and fair ID accuracy evaluation, we propose a new metric for face personalized generation which is noted as ID (Prompt). Specifically, if the CLIP score of this image is lower than the threshold (set as 0.23), then the ID (Prompt) score of this image is 0. The threshold is the average CLIP score of these images which get higher scores in user study. (2) User Study: We select more than $\sim$20 volunteers and generate $\sim$200 images, to evaluate different methods from prompt-image consistency (Prompt), ID accuracy (ID), and image quality (Quality). (3) Parameter Amount: We compare the parameter amount from parameters to be learned (Train) and the total introduced parameters (Add). (4) Time: We evaluate the fine-tuning time.
Figure 5: The different effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens.
Figure 6: Different options for attention loss calculation.
4.2 Comparisons with State-of-the-Art Methods
Qualitative Evaluation. We utilize the same prompt to test five different state-of-the-art personalized generation methods and ours, the qualitative results are shown in Fig. 3. We evaluate the performance from three different levels: facial attributes (e.g., age and hairstyle), actions (i.e., human motion and interactions with other objects), and combinations of facial attributes and actions. Textural Inverion [8] and Prospect [48] tend to overfit the input image, so they fail to interaction with other concepts. Although DreamBooth [30], Custom Diffusion [19] successfully generate the interaction image of human and concept, the generated identities fail to maintain the consistency with the target images. Celeb Basis [47] successfully generate the human-object interaction actions, but they fail to manipulate the facial attributes of target identities well.
@@ 4.2 Comparisons with State-of-the-Art Methods
Table 2: Quantitative evaluation of ablation study.
@@
Quantitative Evaluation. As shown in Tab. 1, our method achieves the SOTA performance in the image-prompt alignment evaluation, and achieves similar face detection rate to Textural Inversion [8]. Due to attention overfit, although Textural Inversion [8]’s ID and Detect evaluation scores are higher than the other methods, it fails to handle the interactive generation between new ID embedding and existing concepts in the original model. For fair ID evaluation, we propose a new metric, which is noted as ID (Prompt), which requires the generated images have to achieve the task of semantic-fidelity, then we calculate their ID scores. Our method achieve better ID (Prompt) scores than Textural Inversion [8] and ours is excellent in image-prompt alignment evaluation, and this improvement is from two reasons: (1) Attention Overfit Alleviation: our face-wise attention loss is able to alleviate the attention overfit problem of previous methods such as DreamBooth [30], Prospect [48], and Texutral Inversion [8]. Our method can make the ID embedding focus on the face region, instead of the whole image. (2) Attribute-Aware Tokens: Compared to Celeb Basis [47], our method does not introduce too much face prior and represent one ID as five per-stage tokens, which can balance the trade-off between ID accuracy and manipulation ability. Our expended textural conditioning space have a strong disentanglement and control ability of attributes (e.g., action-related objects and facial attributes) than Celeb Basis [47].
Figure 7: Utilizing different number of K-V pairs.
User Study. To make our results more convincing, we further conduct a user study, which can be found in Tab. 1. Our method obtains better preference than previous works among the participating users.
Figure 8: Multi-ID embedding and generation.
@@
Efficiency Evaluation: Parameter Amount and Time. As shown in Tab. 1, we have advantages in introduced parameter amount and fine-tuning time. The Celeb Basis [47] introduces a basis module and a pre-trained face recognition model, but these are big optimization burdens and a too strong facial prior can disrupt the interaction between faces and other concepts. We utilize the prior from the T2I model itself to embed IDs, reducing the introduction of additional parameters and further enhancing the facial manipulation ability of the T2I model.
4.3 Ablation Study
@@
Different Effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ Tokens. The semantic information of per-stage tokens is important for the interpretation of diffusion-based generation process, especially for the different effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens. As shown in Fig. 5, we thoroughly explore this from two perspectives: (1) Progressively Adding: We add different ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$ tokens to the conditioning information in ten steps. We found that the initial tokens effect more the layout of generation content (e.g., face region location, and poses), while the latter tokens effect more the ID-related details. (2) Progressively Substituting: We then substitute different $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens of ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$. We found that $\bm{P_{i}^{V}}$ contribute to the vast majority of ID-related conditioning information, and the $\bm{P_{i}^{K}}$ contribute more to textural details, such as environment lighting.
Figure 9: Using our ID embedding method for non-face concepts.
@@
Attention Loss. We thoroughly investigate three options for face-wise attention loss. The option #$\#1$ only regularizes on the $*V*$ token and the option #$\#2$ regularizes the prompt-length tokens. As shown in Fig. 6, option #$\#1$ effects the other concept embeddings in the T2I model, which results in non-ID concepts that cannot be generated, such as sunglasses. Although the option #$\#2$ can reduce the influence of too much ID attention, the activation region of $*V*$ still disrupts regions beyond its scope, which can be seen in the corners of the feature activation map for $*V*$. Our final adopted option is #$\#3$, which calculates the attention loss among the whole text attention maps generated by each token. This option prevents the learned token from overfitting to other regions and only focus on the face region.
@@
The Number of K-V Pairs. As shown in Fig. 7, we explore the influence of K-V pair numbers. When using only one pair of K-V, the learned ID-related tokens fail to maintain good ID accuracy and interact with other complex concepts and attributes. However, adopting too many K-V pairs (e.g., 10 pairs) fails to bring significant improvements of diversity or quality, and this is no doubt a huge computational burden. In our method, we select 5 K-V pairs, which balances the trade-off of representing capacity and computation.
4.4 Generalization: Multi-ID and Other Objects

### 22 Mar 2024 v2

@@ III-C Semantic-Fidelity Token Optimization
Lack of Semantic-Fidelity Control. This problem can be found from two perspectives: (1) Stable Diffusion Model: We observe that even though the face data of celebrities has been included in the training dataset of Stable Diffusion Model, it still fails to achieve perfect semantic-fidelity control for these IDs. For example, “a photo of an old Obama” cannot generate the corresponding images. (2) Previous Personalized Methods: Methods like Celeb Basis [3], Textural Inversion [1] and InstantID [57] mainly emphasize how to preserve the characteristics of the person and achieve global control over the generated images through text modifications. Although these methods are able to manipulate scenes or styles, they struggle to control fine-grained facial attributes of learned IDs, such as age and expressions. Prospect [2] represents an image as a collection of textual token embeddings which could offer better disentanglement and controllability in editing images. However, When it comes to the generation of images with controllable facial attributes, as shown in Fig. 7, it fails to generate examples like “an old person”. We address this challenge by disentangling the $\bm{K}$ and $\bm{V}$ features, as explained in detail in the following section.
@@
Disentanglement of $\bm{K}$ and $\bm{V}$ Features. The text condition features ($\bm{K}$, $\bm{V}$) will be fed into cross-attention layers of U-Nets for conditioning the generated images. Previous methods [8, 50] differentiated the $\bm{K}$ and $\bm{V}$ features calculated from the same ID embedding $\bm{P}$ as position information and object texture features, which is not appropriate for manipulating facial attributes. Therefore, to further investigate the different effects of $\bm{K}$, $\bm{V}$ features for our task, we disentangle the ID embedding $\bm{P}$ as per-stage token embedding groups $\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}_{1\leq i\leq 5}$, and then visualize the effects of these features in the image generation process. As shown in Fig. 4, we found that the $\bm{P^{V}_{i}}$ embeddings are more ID-related, while the $\bm{P^{K}_{i}}$ embeddings are more related to environment factors such as lighting, mouth open and face texture. The disentangled optimization of ID embedding in $\bm{P^{K}_{i}}$ and $\bm{P^{V}_{i}}$ can further improve the ID accuracy and interactive generative ability with other concepts. As shown in Fig. 3, we illustrate the different implementations.
@@
How to Obtain K-V Feature Pairs? Specifically, the input prompt is firstly fed into the CLIP Tokenizer, which generates the textual token embeddings $\bm{Y_{t}}$. Here, the ID embedding $\bm{P}$ related to “$V^{*}$” is a vector with the size of $1\times 768$. As depicted in Fig. 5, the $\bm{P}$ is then fed to a trainable Self-Attention [60] module to create $10\times 768$ embedding $\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}_{1\leq i\leq 5}$. Each of the newly generated ID embedding will replace the original ID embedding to form five groups of textual embeddings, and then these embedding groups will be multiplied by $\bm{W^{k}}$ and $\bm{W^{v}}$ to obtain $\bm{K}$ and $\bm{V}$ features. The Self-Attention module consists of two self-attention layers with one feed-forward layer. We take each group of textual embeddings as a different condition. We evenly divide the 1000 diffusion steps into five stages, each stage corresponds to a unique pair of textual embeddings. Finally, only the Self-Attention module is trainable and the final ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$ is obtained by optimizing the diffusion denoising loss as follows:
$\mathcal{L}_{K,V}=\mathbb{E}_{\bm{z},t,\bm{P}}[||\epsilon-\epsilon_{\theta}(\bm{z_{t}},t,(\bm{P_{i}^{K}},\bm{P_{i}^{V}}))||_{2}^{2}],$ (5)
@@
where $\bm{z}$ is the latent code of target image. The $\epsilon$ is the unscaled noise sample, and the $\epsilon_{\theta}$ is the U-Net module in diffusion model. The $t$ is the optimization step of the diffusion process, and the $\bm{z}_{t}$ is U-Net output of different steps.
Figure 5: The details of Self-Attention module. For simplicity, we disregard the remaining embeddings in $\bm{Y_{t}}$ and focus on the ID embedding $\bm{P}$ associated with the pseudo-word “V*”.
@@
Training Loss. Our goal is seamlessly embedding one specific ID into the space of Stable Diffusion Model, which have to achieve accurate ID mapping and fully use the prior from the Stable Diffusion Model to manipulate scenes, actions and facial attributes. Thus, the total optimization objective can be formulated as follows:
@@
Figure 6: Face photo generation of ours and comparison methods. Due to attention overfitting, Textural Inversion [1] and Prospect [2] struggle to generate images that accurately reflect the semantics of “white hair“. Custom Diffusion [8] and DreamBooth [9] tend to overly mimic the training image and fail to maintain identity when combined with other text prompts. On the other hand, methods such as Celeb Basis [3] and FastComposer [58] exhibit poor semantic fidelity and limited diversity in their generated outputs.
Figure 7: Qualitative comparisons with different SOTA methods using more complex prompts. We conduct experiments from three levels, including the action manipulation, facial attribute editing, and their mixture. Our method shows superior embedding accuracy and interactive generation ability with existing concepts.
IV Experiments
IV-A Experimental Settings
Implementation Details. We present our target T2I model, test data, training details, and inference recipe for reproductivity. (1) Target T2I Model: Unless otherwise specified, we utilize Stable Diffusion 1.4 [4] with default hyper parameters as the pre-trained diffusion-based T2I model. We adopt a frozen CLIP model [33] in the Stable Diffusion Model as the text encoder network. The texts are tokenized into start-token, end-token, and 75 non-text padding tokens. (2) Test Data: The test face images are the StyleGAN [63] synthetic data and the images from the CelebA-HQ dataset [64]. (3) Training Details: For our method, the time for fine-tuning every ID using only one face image is $\sim 5$ minutes ($\sim 1000$ epochs) on one NVIDIA TITAN RTX GPU. We adopt Adam optimizer and set its learning rate as 0.005. The $\lambda$ is set to 0.003. Since we only rely on single face image to acquire its embedding, we adopt some image augmentation methods, including color jitter, horizontal flip with the probability of $0.5$, and random scaling ranging in $0.1\sim 1.0$. (4) Inference Recipe: During sampling time, we employ a DDIM sampler [65] with diffusion steps $T=50$ and the classifier-guidance [66] with the guidance scale $w=7.5$.
Figure 8: The manipulation diversity of our method, which is shown from various identities, styles, facial attributes, and actions.
TABLE I: Quantitative evaluation between different SOTA methods and ours.
@@ IV-A Experimental Settings
Prompt *ID (F) *Detect (F) ID (P) Prompt (U) ID (U) Quality Train Add (min)
@@
TABLE II: Quantitative evaluation using previous metrics.
Prompt ID Detect ID (P)
@@
Baseline Methods. Our task setting is using only one face image to embed the novel ID into the pre-trained Stable Diffusion Model. Thus, for fair comparisons, we only use a single image for all personalized generation methods, but using enough optimization time for different methods. We select six state-of-the-art works as baseline methods for comparisons from three perspectives: (1) Model Fine-Tuning: DreamBooth [9] (learns a unique identifier and fine-tunes the diffusion model to learn from a set of images) and Custom Diffusion [8] (retrieves images with similar captions of the target concept and optimizes the cross-attention module with a modifier token); (2) Token Optimization: Textual Inversion [1] (learns a pseudo-word for a concept within a limited number of images for optimization), ProSpect [2] (expands the textual conditioning space with several per-stage textual token embeddings), and Celeb Basis [3] (builds a well-defined face basis module to constrict the face manifold); (3) Tuning Free: FastComposer [58] (deploys a delayed subject conditioning strategy to achieve tuning-free image generation).
Metrics. We evaluate all the methods from objective metrics, user study, parameter amount, and fine-tuning time. (1) Objective Metrics: We select Prompt (CLIP alignment score [33] between text and image), ID (ID feature similarity score [67]), and Detect (face detection rate [67]). However, evaluating the ID without the essence of T2I generation (i.e., Prompt-Image alignment has the highest priority) is inappropriate, and we DISCUSS the reasons for this problem in Sec. IV-B. Thus, we propose a new metric for face personalized generation which is denoted as ID (P). Specifically, if the CLIP score of this image is lower than the threshold (set as 0.23), then the ID (P) score of this image is 0. The threshold is the average CLIP score of these images which get higher scores in user study. To distinguish these ID metrics, we denote *ID (F) and *Detect (F) for evaluating the images using “a photo of the face of V*”. (2) User Study: We select more than $\sim$20 volunteers and generate $\sim$200 images, to evaluate different methods from Prompt (U) (Prompt-Image alignment), ID (U) (ID accuracy), and Quality (image quality). (3) Parameter Amount: We compare the parameter amount from parameters to be learned (Train) and the total introduced parameters (Add). (4) Time: We evaluate the fine-tuning time of different methods to show efficiency performance.
Figure 9: Multi-ID action manipulation comparisons of Celeb Basis [3], FastComposer [58], and ours. FastComposer only focuses on faces and fails to interact with other concepts, such as “shake hands”, “sofa”, and “picnic”. Although Celeb Basis can generate text-aligned images, it shows lower identity preservation.
Figure 10: Multi-ID scene manipulation comparisons of Celeb Basis [3], FastComposer [58], and ours. As for FastComposer, the faces of target IDs take over most of the generated picture and some concepts are lost, like “garage”. As for Celeb Basis, its learned IDs are less precise and may generate artifacts (i.e., a head of a woman which should not exist in the photo).
Figure 11: Our multi-ID generation results tested in more complex scenarios, showcasing the diversity of generated images and the ability to interact with complex concepts.
IV-B *DISCUSSION* for ID Similarity Evaluation.
As shown in Fig. 6, different from face image generation (i.e., using “a photo of the face of V*”) and editing, achieving Prompt-Image alignment has the highest priority in our task. We have to note this important issue from two perspectives: (1) Explanations for the Previous ID Similarity Metric: As shown in Tab. II, the reason for ID similarity metric less than 0.4, is due to differences in face region resolution. These T2I generated images require face cropping, scaling, and alignment. Consequently, the ID scores are lower than in previous face image generation methods. To fairly evaluate ID similarity under the setting of face photo generation, we conduct ID similarity evaluation using the same-resolution generated face images as the input images with metrics *ID (F) and *Detect (F), as shown in Tab. I. (2) Evaluating ID Considering Text-to-Image Alignment: ID evaluation ignoring the T2I alignment shows “fake high” ID scores. As shown in Fig. 1 and Tab. II, we observe that the previous methods failed to generate the images aligned with prompt such as “a V* is enjoying a cup of latte” and only generate face images due to attention overfitting, but they had the higher ID scores. This ID evaluation ignores pre-requisite of T2I alignment. Therefore, considering the essence of T2I generation, we have to propose ID (P) for fair comparisons.
Figure 12: Ablation study of using per-stage tokens with previous methods. The per-stage tokens strategy enables our method to manipulate the facial attributes of target face, and also works for previous methods.
IV-C Single ID Embedding and Manipulation
We first utilize the same prompts to evaluate five different state-of-the-art methods and ours on single ID embedding and manipulation, as shown in Fig. 7. We evaluate the performance from three different levels: facial attributes (e.g., age and hairstyle), actions (i.e., human motion and interactions with other objects), and combinations of facial attributes and actions. Textual Inversion [1] and Prospect [2] tend to overfit the input image, so they fail to interact with other concepts. Although DreamBooth [9] and Custom Diffusion [8] successfully generate the image of interaction of human and concept, the generated identities fail to maintain the ID consistency with the target images. Celeb Basis [3] successfully generate the human-object interaction actions, but they fail to manipulate the facial attributes of target identities well. Additional results showcased in Fig. 8 further illustrate the diverse range of manipulations accomplished by our methods in terms of scene (stylization), facial attributes, and action representation within the context of single-person image generation.
IV-D Quantitative Evaluation
@@
As shown in Tab. I, our method achieves the SOTA performance in the Prompt-Image alignment evaluation and ID (Face) similarity. Due to attention overfit, Textual Inversion [1], Prospect [2] show poor Prompt-Image alignment. Since that achieving Prompt-Image alignment has the highest priority, we propose a new metric ID (P), which requires the generated images have to achieve the task of semantic-fidelity, and then we calculate their ID scores. Our method achieves better ID (P) scores than the other methods and ours is excellent in Prompt-Image alignment evaluation. This improvement is from two reasons: (1) Attention Overfit Alleviation: our face-wise attention loss is able to alleviate the attention overfit problem of previous methods such as DreamBooth [9], Prospect [2], and Texutal Inversion [1]. Our method can make the ID embedding focus on the face region, instead of the whole image. (2) Attribute-Aware Tokens: Compared to Celeb Basis [3], our method does not introduce too much face prior and represents one ID as five per-stage tokens, which can balance the trade-off between ID accuracy and manipulation ability. Our expended textual conditioning space has a strong disentanglement and control ability of attributes (e.g., action-related objects and facial attributes) than Celeb Basis [3].
Figure 13: Ablation study of different options for attention loss. Option #$\#1$ inferences other concepts, and option #$\#2$ still disrupts regions like the corners of the activation map for $V^{*}$ that beyond its scope.
IV-E Multi-ID Embedding and Manipulation
As shown in Fig. 9 and Fig. 10, we illustrate the circumstances where two IDs appear in the same scene and some interactive actions between them. Though Celeb Basis [3] can achieve competitive prompt alignment as ours, the generated identity is less precise which leads to their poor identity similarity as shown in Tab. I. We hypothesis that in the absence of explicit regularization, the learned ID embedding may be sub-optimal, as they still can not disentangle the identity representation from the other latent factors. For instance, the results in Fig. 9 suggest that their learned ID embedding not only focuses on identity but also incorporates additional information, such as clothing (e.g., the consistent presence of a suited man). In the experiments compared with FastComposer [58], The generated images by FastComposer predominantly feature the faces of the target IDs, occupying a significant portion of the images and it seems like that the characters are directly pasted into the picture, resulting in a disharmonious appearance. Besides, it is difficult for FastComposer to interact with other concepts (like “picnic” and “garage”) and generate the correct action (like “sitting”, “shaking”, and “cooking”) because of the aforementioned semantic prior forgetting problem. As shown in Fig. 11, we experiment on more complex scenarios in multi-ID generation, which showcases the high generation diversity and good interactive ability of our method.
Figure 14: Ablation study of utilizing different number of K-V pairs. Using only 1 K-V pair can not sufficient maintain the ID features. And adopting too many K-V pairs would not bring significant improvements to generation quality. Thus, we finally select 5 K-V pairs.
IV-F User Study
To make our results more convincing and incorporate a broader range of user perspectives, we further conduct a user study, which can be found in Tab. I. Our method obtains better preference than previous work among the participating users, including better Prompt-Image alignment, ID similarity to the target reference image, and image quality. This shows that our semantic-fidelity embedding can enable better interactive generation ability and is potential to exploit the powerful manipulation capabilities of the Stable Diffusion Model itself.
IV-G Efficiency Evaluation
@@
As shown in Tab. I, we have advantages in introduced parameter amount and fine-tuning time. Celeb Basis [3] introduces a basis module and a pre-trained face recognition model, but these are large optimization burdens and a too strong facial prior can disrupt the interaction between faces and other concepts. We utilize the prior from the T2I model itself, reducing the introduction of additional parameters and further enhancing the facial manipulation ability of T2I models.
TABLE III: Quantitative evaluation of ablation study.
[…]
IV-H Ablation Study
@@
Different Effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ Tokens. The semantic information of per-stage tokens is important for the interpretation of diffusion-based generation process, especially for the different effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens. As shown in Fig. 12 we add experiments of using the Per-Stage Token Optimization with previous Textual Inversion, which shows its fine-grained control ability, such as the manipulation of facial attributes. To further investigate this, as shown in Fig. 4, we thoroughly explore $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens from two perspectives: (1) Progressively Adding: We add different ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$ tokens to the conditioning information in ten steps. We found that the initial tokens influence more the layout of generation content (e.g., face region location, and poses), while the latter tokens effect more the ID-related details. (2) Progressively Substituting: We then substitute different $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens of ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$. We found that $\bm{P_{i}^{V}}$ contribute to the vast majority of ID-related information, and the $\bm{P_{i}^{K}}$ contribute more to textual details, such as environment lighting.
Figure 15: Using our ID embedding method for non-face concepts. In each block of part (b), the target object is displayed on the left, while on the right, from top to bottom, are the images labeled as “a photo of V*”, “stylization of V*”, and “V* under different scenes”.
Figure 16: Using our ID embedding method for Stable Diffusion XL. InstantID [57] tends to generate a face photo of target ID. PhotoMaker [59] and IP-Adapter-FaceID [68] can not achieve fine-grained text guided facial attribute controlling. Our method can achieve better interactive generation with the other concepts (e.g., chair) than the other methods.
@@
Attention Loss. We thoroughly investigate three options for face-wise attention loss. The option #$\#1$ only regularizes on the $V^{*}$ token and the option #$\#2$ regularizes the prompt-length tokens. As shown in Fig. 13, option #$\#1$ affects the other concept embeddings in the T2I model, which results in non-ID concepts cannot be generated, such as sunglasses. Although the option #$\#2$ can reduce the influence of too much ID attention, the activation region of $V^{*}$ still disrupts regions beyond its scope, which can be seen in the corners of the feature activation map for $V^{*}$. Our final adopted option is #$\#3$, which calculates the attention loss among the whole text attention maps generated by each token. This option prevents the learned token from overfitting to other regions and only focus on the face region. Drawn from Tab. III, as more tokens are token into the attention loss regularization, the prompt score rises. We think the reason lies in two perspective: (1) The regularization on the $V^{*}$ token ensures it to focus on face region and prevents it from disturb the other concepts; (2) The regularization applied to all other tokens serves as an additional penalty, preserving their ability to implicitly disentangle the $V^{*}$ token from the rest of the tokens. Our loss strategy only addresses the attention overfitting, improving the ID accuracy and interactivity with other concepts, but the manipulation capacity for the high text2image alignment and diversity still needs to be improved by us and other diffusion-based generative model researchers.
@@
The Number of K-V Feature Pairs. As shown in Fig. 14, we explore the influence of K-V pair numbers. When using only one pair of K-V, the learned ID-related tokens fail to maintain good ID accuracy and interact with other complex concepts and attributes. However, adopting too many K-V pairs (e.g., 10 pairs) fails to bring significant improvements of diversity or quality, and this is no doubt a huge computational burden. In our method, we select 5 K-V pairs, which balance the trade-off of representing capacity and computation. As shown in Tab. III, the Prompt and identity scores of setting 1 K-V with option #$\#3$ are lower than 5 K-V with option #$\#3$ and 10 K-V with option #$\#3$. While the 10 K-V with option #$\#3$ shows the same prompt score compared to 5 K-V with option #$\#3$, it exhibits lower identity similarity.
IV-I Generalization